A 61.26 gram sample of iron (with a specific heat of 0.450 J/goC) is heated to 100.0 ºC. It is then transferred to a coffee cup calorimeter containing 49.61 g of water (specific heat of 4.184 J/ g ºC) initially at 20.63 ºC. If the final temperature of the system is 23.59, what was the heat absorbed (q) of the calorimeter? (Hint: remember that the total heat absorbed by the water AND the calorimeter = the heat released by the iron.) Record your answer as a whole number (assume the sign is positive).
1
Expert's answer
2019-04-23T01:59:08-0400

QFe = QH2O + Qcal.

QFe = CFe × mFe × (T2Fe – T1Fe).

QH2O = CH2O × mH2O × (T2H2O – T1H2O).

0.450 J/goC × 61.26 g × (100.0 oC – 23.59 oC) = 4.184 J/goC × 49.61 g × (23.59 oC – 20.63 oC) + Qcal.

2106.3945 J = 614.4020 J + Qcal.

Qcal = 2106.3945 J – 614.4020 J = 1491.9925 J ≈ 1492 J.